# DynaSeg: A Deep Dynamic Fusion Method for Unsupervised Image Segmentation Incorporating Feature Similarity and Spatial Continuity

### v3

@@ 5.3 Comparison Between FSF and SCF
To further refine the comparative analysis between the Feature Similarity Focus (FSF) and Spatial Continuity Focus (SCF) methods, we consider additional mIOU scores on BSD500 and PASCAL VOC2012 datasets, as presented in Table 1.
In the BSD500 dataset, across different segmentation scenarios (All, Fine, and Coarse), FSF consistently outperforms SCF in terms of mIOU scores. Notably, for BSD500 All, FSF achieves an mIOU of 0.349 compared to SCF’s 0.330. This pattern is observed in other scenarios, reinforcing the effectiveness of FSF in capturing fine-grained details and semantic nuances.
On the PASCAL VOC2012 dataset, both FSF and SCF demonstrate competitive performance, with FSF achieving an mIOU of 0.391 and SCF’s 0.396. The marginal difference in mIOU scores on PASCAL VOC2012 suggests that both methods perform comparably on this specific dataset, indicating their capability to handle diverse segmentation challenges effectively.
Comparing the two methods across different datasets and segmentation scenarios, FSF consistently delivers higher mIoU scores, indicating its robustness and effectiveness in capturing both global and fine-grained semantic information. The choice of FSF over SCF is particularly beneficial when targeting scenarios with varied and intricate segmentation requirements.
Both FSF and SCF perform better with a ResNet-18 backbone compared to CNN-based backbones. FSF, in particular, achieves higher mIoU values for ”Stuff” across both backbones, emphasizing its effectiveness in capturing detailed and intricate features. Therefore, FSF shows promise for scenarios involving complex and diverse categories, especially when coupled with a ResNet-18 backbone.

### v4

@@ 5 Ablation Study
5.1 Silhouette Score Impact
The accompanying graph in Figure 6 provides additional insights into the distribution of ground truth cluster numbers, complementing the tabular results. The graph illustrates a significant disparity when comparing the distribution with a random threshold of 3. Conversely, when comparing the distribution with the optimal number of clusters derived from the Silhouette Score, a strong correlation is observed. This finding underscores the effectiveness of the proposed dynamic weighting scheme.
5.2 Effect of Feature Extractor
@@ 5.3 Comparison Between FSF and SCF
To further refine the comparative analysis between the Feature Similarity Focus (FSF) and Spatial Continuity Focus (SCF) methods, we consider additional mIOU scores on BSD500 and PASCAL VOC2012 datasets, as presented in Table 1. In the BSD500 dataset, across different segmentation scenarios (All, Fine, and Coarse), FSF consistently outperforms SCF in terms of mIOU scores. Notably, for BSD500 All, FSF achieves an mIOU of 0.349 compared to SCF’s 0.330. This pattern is observed in other scenarios, reinforcing the effectiveness of FSF in capturing fine-grained details and semantic nuances. On the PASCAL VOC2012 dataset, both FSF and SCF demonstrate competitive performance, with FSF achieving an mIOU of 0.391 and SCF’s 0.396. The marginal difference in mIOU scores on PASCAL VOC2012 suggests that both methods perform comparably on this specific dataset, indicating their capability to handle diverse segmentation challenges effectively.
Comparing the two methods across different datasets and segmentation scenarios, FSF consistently delivers higher mIoU scores, indicating its robustness and effectiveness in capturing both global and fine-grained semantic information. The choice of FSF over SCF is particularly beneficial when targeting scenarios with varied and intricate segmentation requirements. Both FSF and SCF perform better with a ResNet-18 backbone compared to CNN-based backbones. FSF, in particular, achieves higher mIoU values for ”Stuff” across both backbones, emphasizing its effectiveness in capturing detailed and intricate features. Therefore, FSF shows promise for scenarios involving complex and diverse categories, especially when coupled with a ResNet-18 backbone.